# Fundamental limitations on the recoverability of quantum processes

### v1

@@ Acknowledgments
VP and Sohail thank the International Institute of Information Technology Hyderabad for the support and hospitality during their visits in the summer 2023 where part of this work was done. SD acknowledges support from the Science and Engineering Research Board, Department of Science and Technology (SERB-DST), Government of India under Grant No. SRG/2023/000217. SD also thanks the International Institute of Information Technology Hyderabad for the Faculty Seed Grant.
@@ Appendix A Structure of the Appendix
We discuss detailed proofs of the meta and main results (lemmas, propositions, and theorems) and some additional observations in the Appendix. The structure of the Appendix is as follows. In Appendix B, we provide the proofs of statements (which are about some desirable properties of spaces of channels and superchannels) from Section II. In Appendix C, we have proof of statements from Section III. This section also contains sesevral bounds on entropy gain under CP, CP unital, and positive maps ( see Proposition 13, Lemma 10 and Theorem 3). In Appendix D, we provide the proof of the Theorem 2, which is also one of our main results and it provides a refinement of the data processing inequality. In Appendix E we provide the proof of monotonicity of $\mathbf{D}^{(n)}$.
@@ B.2 Proof of Lemma 2
We will first show that “if part" holds. It can be shown that $\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)$ is CP. As we have assumed that the supermap $\Theta$ is completely CP-preserving, we have $\Lambda_{\Theta}=\left(\mathrm{id}\otimes\Theta\right)\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)$ is CP.
We will now prove the nontrivial “only if" part. Given that $\Lambda_{\Theta}$ is CP, we need to prove that if $\Omega\in\mathcal{L}\left(E,E^{\prime}\right)\otimes\mathcal{L}(A,B)$ is CP (where $E$ and $E^{\prime}$ are reference systems), then $\left({\mathrm{id}}\otimes\Theta\right)\Omega$ is also CP. First, we prove the following lemma
@@
$\Omega=\sum_{n}\sum_{\{\alpha,\beta,\gamma,\delta\}}\sum_{\{p,q,r,s\}}{\kappa}^{n}_{\alpha\beta\gamma\delta}\bar{{\kappa}}^{n}_{pqrs}\mathcal{E}^{\mathcal{K}_{E}\rightarrow\mathcal{K}_{E^{\prime}}}_{\gamma r\alpha p}\otimes\mathcal{E}^{A\rightarrow B}_{\delta s\beta q},$ (89)
where $\kappa^{n}_{\alpha\beta\gamma\delta}$ are complex numbers, $\bar{\kappa}^{n}_{pqrs}$ denotes the complex conjugate of $\kappa^{n}_{pqrs}$ (these complex numbers depends on the CP map $\Omega$.), and $\{\mathcal{E}^{\mathcal{K}_{E}\rightarrow\mathcal{K}_{E^{\prime}}}_{\gamma r\alpha p}\}_{\gamma r\alpha p}$, $\{\mathcal{E}^{A\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ are orthonormal basis for $\mathcal{L}\left(E,E^{\prime}\right)$ and $\mathcal{L}\left(A,B\right)$, respectively.
@@
Let $\{e^{Q}_{ij}\}_{ij}$, where $Q=\{A,B,E,E^{\prime}\}$ denotes a canonical basis for the spaces $\mathcal{L}\left({A}\right)$, $\mathcal{L}\left({B}\right)$, $\mathcal{L}\left({E}\right)$, and $\mathcal{L}\left({E^{\prime}}\right)$, respectively. In the bra-ket notation we can write $e^{Q}_{ij}=\outerproduct{i_{Q}}{j_{Q}}$ |, where {|$\{\ket{i_{Q}}\}_{i}$ is an orthonormal basis for $Q$. A map $\Omega$ is CP if and only if it has a Choi-Kraus decomposition as follows
@@
The second line simply follows from the definition of the Hilbert-Schmidt inner product and we have written back everything on a canonical basis. In the third line, we have used a decomposition of $\mathsf{X}=\sum_{\{ijkl\}}x_{ijkl}e^{E}_{ij}\otimes e^{A}_{kl}$ to rewrite $\mathsf{X}$ in terms of canonical basis, and $\{\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\}_{\gamma r\alpha p}$, $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ are basis for spaces $\mathcal{L}\left(E,E^{\prime}\right)$ and $\mathcal{L}\left(A,B\right)$, respectively. The fourth and fifth line follows from the definition of basis in $\{\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\}$ and $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}$ (see Eq.(17)). This concludes the proof of the lemma. ∎
@@
where $\kappa^{n}_{\alpha\beta\gamma\delta}$ are complex numbers defined in Eq. (91) and $\bar{\kappa}^{n}_{\alpha\beta\gamma\delta}$ are complex conjugate of $\kappa^{n}_{\alpha\beta\gamma\delta}$. So, we can write $\Omega$ as follows
@@
It was shown in Lemma (8) that a map $\Lambda_{\Theta}\in\mathcal{L}(A,B)\otimes\mathcal{L}(C,D)$ is CP if and only if it can be written as
@@ B.3 Proof of Lemma 3
We already have shown a supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is completely CP-preserving if and only if $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)\in\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)$ is CP, where $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ are basis for $\mathcal{L}\left(A,B\right)$. We can determine the complete positivity of $\Lambda_{\Theta}$ by checking the positivity of its Choi operator $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ as defined below
@@
where $\{e^{Q}_{ij}\}_{ij}$ denotes a canonical basis for the spaces $\mathcal{L}\left(Q\right)$ for $Q=\{A,B,C,D\}$. Using the form of $\Lambda_{\Theta}$ from Eq. (18) in the above equation, we have
@@
Now, the transformation $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}\leftrightarrow\widehat{\mathsf{C}}_{\mathsf{T}}$ is an ${}^{\ast}$-isomorphism. So, $\widehat{\mathsf{C}}_{\mathsf{T}}$ is positive semidefinite if and only if $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ is positive positive semidefinite. Now, $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ is positive semidefinite if and only if $\Lambda_{\Theta}$ is CP, and $\Lambda_{\Theta}$ is CP if and only if $\Theta$ is completely CP-preserving. As $\mathsf{T}$ is CP if and only if $\widehat{\mathsf{C}}_{\mathsf{T}}$ is positive semidefinite, we have $\mathsf{T}$ is CP if and only if $\Theta$ is completely CP-preserving. This completes the proof.
B.4 Proof of Lemma 4

### v2

@@ Acknowledgments
We are grateful to Maarten Wegewijs and Kaiyuan Ji for carefully reading our paper and pointing out an error to a previous version of Theorem 2. We thank Mark M Wilde for discussions. VP and Sohail thank the International Institute of Information Technology Hyderabad for the support and hospitality during their visits in the summer 2023 where part of this work was done. SD acknowledges support from the Science and Engineering Research Board, Department of Science and Technology (SERB-DST), Government of India under Grant No. SRG/2023/000217. US and SD thank the International Institute of Information Technology Hyderabad for the Faculty Seed Grant.
@@ Appendix A Structure of the Appendix
In the appendices, we provide detailed proofs of the results (lemmas, propositions, and theorems) of the main text and offer some additional observations. The structure of the appendix is as follows. In Appendix B, we provide the proofs of statements (which are about some desirable properties of spaces of channels and superchannels) from Section II. In Appendix C, we have proof of statements from Section III. This section also contains several bounds on entropy gain under CP, CP unital, and positive maps (see Theorem 3, Lemma 10 and Proposition 13). In Appendix D, we provide the proof of Theorem 2, which is also one of our main results and it provides a refinement of the data processing inequality. In Appendix E, we provide the proof of monotonicity of $\mathbf{D}^{(n)}$.
@@ B.1 Proof of Lemma 1
$=\sum_{i}\left<\mathcal{N}\left(\mathsf{B}^{A}_{i}\right),\mathcal{M}(a^{A}_{i})\right>$
$=\sum_{i}\left<\mathcal{N}(a^{A}_{i}),\mathcal{M}(a^{A}_{i})\right>,$
where $\{a^{A}_{i}\}$ is some orthonormal basis in $\mathcal{L}\left({A}\right)$ and we have used the fact that $\mathsf{B}^{A}_{i}:=\sum_{j}\left<b^{A}_{j},a^{A}_{i}\right>b^{A}_{j}=a^{A}_{i}$ in the last equality. This concludes the proof.
@@ B.2 Proof of Lemma 2
We will first show that "if part" holds. Note that $\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\mathcal{E}^{A\rightarrow B}_{ijkl}$ is a CP map. As we have assumed that the supermap $\Theta$ is completely CP-preserving, we have $\Lambda_{\Theta}=\left(\mathrm{id}\otimes\Theta\right)\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\mathcal{E}^{A\rightarrow B}_{ijkl}$ is CP.
We will now prove the nontrivial "only if" part. Given that $\Lambda_{\Theta}$ is CP, we need to prove that if $\Omega\in\mathcal{L}\left(E,E^{\prime}\right)\otimes\mathcal{L}(A,B)$ is CP (where $E$ and $E^{\prime}$ are reference systems), then $\left({\mathrm{id}}\otimes\Theta\right)\Omega$ is also CP. First, we prove the following lemma
@@
$\Omega=\sum_{n}\sum_{\{\alpha,\beta,\gamma,\delta\}}\sum_{\{p,q,r,s\}}{\kappa}^{n}_{\alpha\beta\gamma\delta}\bar{{\kappa}}^{n}_{pqrs}\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\otimes\mathcal{E}^{A\rightarrow B}_{\delta s\beta q},$ (89)
where $\kappa^{n}_{\alpha\beta\gamma\delta}$ are complex numbers, $\bar{\kappa}^{n}_{pqrs}$ denotes the complex conjugate of $\kappa^{n}_{pqrs}$ (these complex numbers depend on the CP map $\Omega$), and $\left\{\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\right\}_{\gamma r\alpha p}$, $\left\{\mathcal{E}^{A\rightarrow B}_{\delta s\beta q}\right\}_{\delta s\beta q}$ are orthonormal basis for $\mathcal{L}\left(E,E^{\prime}\right)$ and $\mathcal{L}\left(A,B\right)$, respectively.
@@
Let $\left\{e^{Q}_{ij}\right\}_{ij}$ be a canonical basis for the space $\mathcal{L}\left({Q}\right)$, where $Q=\{A,B,E,E^{\prime}\}$. In the bra-ket notation we can write $e^{Q}_{ij}=\outerproduct{i_{Q}}{j_{Q}}$ |, where {|$\left\{\ket{i_{Q}}\right\}_{i}$ is an orthonormal basis for $Q$. A map $\Omega$ is CP if and only if it has a Choi-Kraus decomposition as follows
@@
The second line simply follows from the definition of the Hilbert-Schmidt inner product and we have written back everything in a canonical basis. In the third line, we have used a decomposition of $\mathsf{X}=\sum_{\{ijkl\}}x_{ijkl}e^{E}_{ij}\otimes e^{A}_{kl}$ to rewrite $\mathsf{X}$ in terms of canonical basis, and $\{\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\}_{\gamma r\alpha p}$, $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ are basis for spaces $\mathcal{L}\left(E,E^{\prime}\right)$ and $\mathcal{L}\left(A,B\right)$, respectively. The fourth and fifth line follows from the definition of basis in $\{\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\}$ and $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}$ (see Eq.(17)). This concludes the proof of the lemma.
@@
where $\kappa^{n}_{\alpha\beta\gamma\delta}$ are complex numbers defined in Eq. (91) and $\bar{\kappa}^{n}_{\alpha\beta\gamma\delta}$ is the complex conjugate of $\kappa^{n}_{\alpha\beta\gamma\delta}$. So, we can write $\Omega$ as follows
@@
It was shown in Lemma 8 that a map $\Lambda_{\Theta}\in\mathcal{L}(A,B)\otimes\mathcal{L}(C,D)$ is CP if and only if it can be written as
@@ B.3 Proof of Lemma 3
We already have shown a supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is completely CP-preserving if and only if $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)\in\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)$ is CP, where $\left\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\right\}_{\delta s\beta q}$ is a basis for $\mathcal{L}\left(A,B\right)$. We can determine the complete positivity of $\Lambda_{\Theta}$ by checking the positivity of its Choi operator $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ as defined below
@@
where $\{e^{Q}_{ij}\}_{ij}$ denotes a canonical basis for the space $\mathcal{L}\left(Q\right)$ and $Q=\{A,B,C,D\}$. Using the form of $\Lambda_{\Theta}$ from Eq. (18) in the above equation, we have
@@
Now, the transformation $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}\leftrightarrow\widehat{\mathsf{C}}_{\mathsf{T}}$ is an -isomorphism. So, $\widehat{\mathsf{C}}_{\mathsf{T}}$ is positive semidefinite if and only if $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ is positive semidefinite. Now, $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ is positive semidefinite if and only if $\Lambda_{\Theta}$ is CP, and $\Lambda_{\Theta}$ is CP if and only if $\Theta$ is completely CP-preserving. As $\mathsf{T}$ is CP if and only if $\widehat{\mathsf{C}}_{\mathsf{T}}$ is positive semidefinite, we have $\mathsf{T}$ is CP if and only if $\Theta$ is completely CP-preserving. This completes the proof.
B.4 Quantum channels from representing maps
Let us define a map $\mathsf{T}^{\prime}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ as follows
$\mathsf{T}{{}^{\prime}}(\mathsf{X})=\mathsf{T}(\mathsf{X})+\mathsf{T}_{0}(\mathsf{X}),$ (99)
where $\mathsf{T}_{0}(\mathsf{X}):=\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0}$, $\sigma_{0}\in\mathcal{L}(C\otimes D)$ and $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ is the representing map of the superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$. Note that the representing map $\mathsf{T}$ is CP but it could be trace non-increasing or trace increasing or neither trace increasing nor trace decreasing. However, if we take $\sigma_{0}$ to be an operator with unit trace then $\mathsf{T}^{\prime}$ becomes a trace preserving map but then $\mathsf{T}^{\prime}$ may not be a CP map. In the following, we wish to find out if $\mathsf{T}^{\prime}$ is a CP map given that $\mathsf{T}$ is CP and $\operatorname{tr}(\sigma_{0})=1$. In the case where $\mathsf{T}$ is trace non-increasing, the condition $\sigma_{0}\geq 0$ in Eq. (99) guarantees that the map $\mathsf{T}^{\prime}$ is CP. In other cases, we derive the necessary and sufficient conditions for $\mathsf{T}^{\prime}$ to be CP as follows.
Necessary and Sufficient Conditions.—The Choi operator of the map $\mathsf{T}_{0}$ in Eq. (99) is given by
$\widehat{\mathsf{C}}_{\mathsf{T}_{0}}=\sum_{ij}e_{ij}\otimes\mathsf{T}_{0}(e_{ij})=\sum_{ij}e_{ij}\otimes\left[\delta_{ij}-\mathrm{tr}(\mathsf{T}(e_{ij}))\right]\sigma_{0}=\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}\otimes\sigma_{0},$ (100)
where $t$ denotes the transpose operation in the basis $\{e_{ij}\}$. Then Eq. (99) in terms of Choi operators implies
$\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}=\widehat{\mathsf{C}}_{\mathsf{T}}+\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}\otimes\sigma_{0},$ (101)
where $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ and $\widehat{\mathsf{C}}_{\mathsf{T}}$ are the Choi operators for the maps $\mathsf{T}^{\prime}$ and $\mathsf{T}$, respectively. We have freedom over $\sigma_{0}$ in the above equation and we need to choose it such that $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ becomes positive semidefinite. Note that $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}$ is a hermitian operator and let its spectral decomposition be given by $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}=\sum_{m:\lambda_{m}>0}\lambda_{m}\Pi_{m}-\sum_{m:\lambda_{m}<0}|\lambda_{m}|\Pi_{m}+\sum_{m:\lambda_{m}=0}\lambda_{m}\Pi_{m}$, where $\{\Pi_{m}\}_{m:\lambda_{m}>0},~{}\{\Pi_{m}\}_{m:\lambda_{m}<0}$ and $\{\Pi_{m}\}_{m:\lambda_{m}=0}$ are orthogonal projectors. Further, define $\Pi_{+}:=\sum_{m:\lambda_{m}>0}\Pi_{m}$ and $\Pi_{-}:=\sum_{m:\lambda_{m}<0}|\lambda_{m}|\Pi_{m}$ as the projectors onto the support of positive and negative eigenvalues of the hermitian operator $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}$, respectively. We denote by $\Pi_{0}$ the projector onto the eigenspace corresponding to the zero eigenvalue of $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}$ One can now prove easily that $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ is positive semidefinite if and only if the operator $\sigma_{0}\in\mathcal{L}(C\otimes D)$ satisfies
$\left(\Pi_{+}-\Pi_{-}\right)\otimes\sigma_{0}\leq\left(\left[|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}]\otimes\mathbbm{1}_{CD}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left([\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}]\otimes\mathbbm{1}_{CD}\right).$ (102)
Now we will show that this condition is also compatible with the trace condition on $\sigma_{0}$ ($\operatorname{tr}\sigma_{0}=1$). By taking the trace with respect to the second subsystem in the above equation, we obtain
$\left(\Pi_{+}-\Pi_{-}\right)(\operatorname{tr}{\sigma_{0}})\leq\left(\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}\right)(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}\left(\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}\right).$ (103)
Let $\{\mu_{k}\}$ be the spectrum of $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}$, then the spectral decomposition of $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}$ gives $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}=\sum_{i}\mu_{i}P_{i}$, where $P_{i}$’s are the projectors onto the eigensubspaces of $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}$. Then we have $\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}=\sum_{i}\left|\mu_{i}-1\right|^{-\frac{1}{2}}P_{i}$, which implies $\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}+\Pi_{0}\right)(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}+\Pi_{0}\right)=\sum_{i}\frac{\mu_{i}}{\left|\mu_{i}-1\right|}P_{i}+\Pi_{0}$. Note that $\mu_{i}$ is positive but can be less than one while the operator $\left(\Pi_{+}-\Pi_{-}\right)$ may be nonpositive, then there is no clash of trace condition on $\sigma_{0}$, i.e., $\operatorname{tr}(\sigma_{0})=1$, with inequality in Eq. (103).
In the special case where $\mathsf{T}$ is a trace increasing map, i.e., $\operatorname{tr}(\mathsf{X})<\mathrm{tr}(\mathsf{T}(\mathsf{X}))$ for all $X\in\mathcal{L}\left(A\otimes{B}\right)$ or $\mathsf{T}^{*}(\mathbbm{1}_{CD})-\mathbbm{1}_{AB}$ is a strictly positive operator. Under this condition we have $\Pi_{+}=\mathbbm{I}_{AB}$ and $\Pi_{-}=0$ and $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ is positive semidefinite if and only if the operator $\sigma_{0}\in\mathcal{L}(C\otimes D)$ satisfies
$\mathbbm{1}_{AB}\otimes\sigma_{0}\leq\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right).$ (104)
Taking partial trace with respect to the subsystems $C$ and $D$ in above equation, we get
$\mathbb{I}_{AB}\operatorname{tr}(\sigma_{0})\leq\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\right)(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\right)=\sum_{i}\frac{\mu_{i}}{\mu_{i}-1}P_{i}>\mathbbm{I}_{AB}.$ (105)
Thus the trace condition on $\sigma_{0}$, i.e., $\operatorname{tr}(\sigma_{0})=1$ does not create any contradiction between Eq. (104).
We will now provide two examples on how to construct $\sigma_{0}$ for a given trace increasing map $\mathsf{T}$ another map $\mathsf{M}$ that is neither trace increasing nor trace decreasing.
Example $1$:–Let us define a map $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ as follows
$\mathsf{T}(\mathsf{X})=\alpha\operatorname{tr}(\mathsf{X})\mathbbm{1}_{CD},$ (106)
where $\alpha$ is real. We are not assuming whether $\mathsf{T}$ is trace increasing or decreasing yet. The above equation also implies $\mathsf{T}^{*}(\mathbbm{1}_{CD})=\alpha n\mathbbm{1}_{AB}$, where $n:=\operatorname{tr}(\mathbbm{1}_{CD})$ and $\mathsf{T}^{*}$ is adjoint of $\mathsf{T}$. We now have
$\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right)=\left(\frac{\alpha}{n\alpha-1}\right)\mathbbm{1}_{AB}\otimes\mathbbm{1}_{CD}.$ (107)
Notice that if we choose $\alpha>\frac{1}{n}$, then $\mathsf{T}$ becomes trace increasing map. Using Eq. (104) we obtain the following inequality
$\mathbbm{1}_{AB}\otimes\sigma_{0}\leq\left(\frac{\alpha}{n\alpha-1}\right)\mathbbm{1}_{AB}\otimes\mathbbm{1}_{CD}.$ (108)
By a suitable choice of $\alpha$, we can make the number $\left(\frac{\alpha}{n\alpha-1}\right)$ as large as possible: take $\alpha=\frac{1}{n}+\frac{1}{n^{2}L}$, where $L$ is a very large positive real number, then $\left(\frac{\alpha}{n\alpha-1}\right)=L+\frac{1}{n}$. If we choose $\sigma_{0}$ to a unit trace and diagonal matrix with the diagonal elements being less than $L+\frac{1}{n}$ (some of the diagonal entries can be chosen to be negative also), then the above inequality will hold. Thus with this choice of $\sigma_{0}$, the CP and trace increasing map $\mathsf{T}(\mathsf{X})=\alpha\operatorname{tr}(\mathsf{X})\mathbbm{1}_{CD}$ can be made a CP and trace preserving map, using Eq. (99).
Example $2$:– Let us define a map $\mathsf{T}$, whose input and output systems are qubits, by the Kraus operators $K_{1}=\alpha\outerproduct{0}{0}$ | and $K_{2}=\beta\outerproduct{1}{1}$ |, where {|$\{\ket{0},\ket{1}\}$ are computational basis and $\alpha>1$, $10<\beta<$ 1 are positive real constants. If we take the input of $\mathsf{T}$ as $\mathsf{X}=\outerproduct{0}{0}$ |, we then have $\operatorname{tr}(\mathsf{X})<\mathrm{tr}(\mathsf{T}(\mathsf{X}))$. On the other hand if we take the input of $\mathsf{T}$ as $\mathsf{X}=\outerproduct{1}{1}$ |, we then have $\operatorname{tr}(\mathsf{X})>\mathrm{tr}(\mathsf{T}(\mathsf{X}))$. Clearly the map $\mathsf{T}$ is neither trace increasing nor trace decreasing map. The Choi operator of $\mathsf{T}$ is given by
$\widehat{\mathsf{C}}_{\mathsf{T}}=\alpha^{2}\outerproduct{00}{00}+\beta^{2}\outerproduct{11}{11}.$ | . (109)
Now we will calculate the right hand side of Eq. (102). First we need the following term
[…]
We then have
$\mathsf{A}:=\left(\left|(\mathsf{T}^{*}(\mathbbm{1}))^{t}-\mathbbm{1}\right|^{-\frac{1}{2}}\otimes\mathbbm{1}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left(\left|(\mathsf{T}^{*}(\mathbbm{1}))^{t}-\mathbbm{1}\right|^{-\frac{1}{2}}\otimes\mathbbm{1}\right)=\alpha^{2}(\alpha^{2}-1)^{-1}\outerproduct{00}{00}+\beta^{2}(\beta^{2}-1)^{-1}\outerproduct{11}{11}.$ | . (111)
We can write $\mathsf{T}^{*}(\mathbbm{1})^{t}-\mathbbm{1}=M-N$, where $M=(\alpha^{2}-1)\outerproduct{0}{0}$ | and $N=(1-\beta^{2})\outerproduct{1}{1}$ |. This also implies $\Pi_{M}=\outerproduct{0}{0}$ | and $\Pi_{N}=\outerproduct{1}{1}$ |. We want to find a $\sigma_{0}$ such that $\operatorname{tr}(\sigma_{0})=1$ and $(\Pi_{M}-\Pi_{N})\otimes\sigma_{0}\leq\mathsf{A}$. Let us choose a diagonal matrix $\sigma$ with eigenvalues $\lambda_{1}$ and $\lambda_{2}$ such that $\lambda_{1}<\alpha^{2}(\alpha^{2}-1)^{-1}$, $\left|\lambda_{2}\right|<\beta^{2}(\beta^{2}-1)^{-1}$, $\lambda_{2}<0$ and $\lambda_{1}+\lambda_{2}>1$. Now if we take $\alpha=\sqrt{2}$, $\beta=\frac{1}{\sqrt{2}}$ and choose $\lambda_{1}=1.99$, $\lambda_{2}=-\frac{1}{5}$ then all the above-mentioned condition are satisfied. Now we choose $\sigma_{0}=\frac{\sigma}{\operatorname{tr}(\sigma)}$. With this choice we have $\operatorname{tr}(\sigma_{0})=1$ and $(\Pi_{M}-\Pi_{N})\otimes\sigma_{0}\leq\mathsf{A}$. This example can trivially be generalized to incorporate bipartite systems suitable in the context of representing maps.
B.5 Proof of Lemma 4